1. Consider the following numerical example of the IS-LM model: C = 100 + 0.3YD
I...

1. Consider the following numerical example of the IS-LM model:

C = 100 + 0.3YD
I = 150 + 0.2Y - 1000i
T = 100
G = 200
i = .01 (M/P)s = 1200
(M/P)d = 2Y - 4000i

a. Find the equation for aggregate demand (Y).

b. Derive the IS relation.

c. Derive the LM relation if the central bank sets an interest
rate of 1%.

d. Solve for the equilibrium values of output, interest rate, C
and I.

e. Expansionary monetary policy. Suppose that the central
bank increases money supply to 1500. What is the impact of
this expansionary monetary policy on the IS and LM curves?
Find the new equilibrium values of output, interest rate, C 
and I.

f. Expansionary fiscal policy. Suppose that the government
increases its spending G to 300. What is the impact of
this expansionary fiscal policy on the IS and LM curves?
Find the new equilibrium values of output, interest rate, C 
and I.
Answer #1

C= 100 +0.3% I= 100 G=200 I. - 150 +0.27 - 1000 iz o.o1 (M/P)s = 1200 (M/Pd = 2Y - 4000 i a Y=c+ Ito disposen e - [ we know Y- 1000 i Ya 100+ 0.3C Y-100) + 150 +0.27 -1000 +200 = 450 +0.3Y- 30 to. 27 - 1000 r=420 +0.57 - 1000 i 0.5 Y = 420 -1000 i (rUy At legem in money mrt (A) a = (4) Money supply = money demand 27-4000i = 1200 24 = 1200 + 4000 - y =600 + 2000 i ₃ LM wineC Cat y = 120) = 100 +0.3 Yo = 100 10:30Y-T) = 100 +0.3( 720 - 100) (* = 286 Eqm consumption is 286 I fat Y = 720) = 150 to.2Week National Day (Saudi Arabia) In IS-LM len - Is= LM is: Y = 840 - 2000 i LM: Y=750 +2000 i 75072000 i 840 - 20000 = 90 - 4Heritage t o 6 expansionary fisial folicy C G is to 300 - 8 =) is curre will shift A the right you nuo is will be: Ye C+ I +GWeek: 37 new interest rate is o ll or 11t. 600 + 2000 (0.11) Year i=0.11) - - 1* = 820 , м 1 и ш 8 umu c cat y=820) = 100 +0.